OpenStudy (anonymous):

what number is the midpoint between two-thirds and 1

OpenStudy (anonymous):

1+2/3= 5/3 mid way is the same as half way. 5/3 divided by 2

OpenStudy (anonymous):

idk what the answer is

OpenStudy (anonymous):

one sec.

OpenStudy (anonymous):



OpenStudy (anonymous):

So the "distance" from 2/3 to 1 = 1/3

OpenStudy (anonymous):

I don't have that option my options are 3/4, 5/6 or 7/9

OpenStudy (anonymous):



OpenStudy (anonymous):

half of 1/3= 1/6

OpenStudy (anonymous):

2/3+1/6= 5/6